A Bibliography of Publications of Ilse C. F. Ipsen

Ilse C. F. Ipsen North Carolina State University

Department of Mathematics

P.O. Box 8205 Raleigh, NC 27695-8205

USA

Tel: +1 919 515 3399 FAX: +1 919 515 3798

E-mail: ipsen@math.ncsu.edu (Internet)

29 August 2024 Version 1.44

Abstract

This bibliography records publications of Ilse C. F. Ipsen.

Title word cross-reference

p [CI21]. QR [CI91, CI94d, CI95a, HIW15, Ips84a]. sin Θ [Ips00c].

-norms [CI21].

14th [IMŠ^{+}09a, IMŠ^{+}09b].

2004 [Ips06f]. 2007 [IMŠ^{+}09a, IMŠ^{+}09b].

’92 [xJ94].

Absolute [EI98b, Ips00b, EI99, Ips03]. Accuracy [BCCI98, JI92]. accurate [Ips06e, Ips06f, Ips09f]. Active [HIS18]. Advanced [GD88]. Algebra [GD88, IMŠ^{+}09b, xJ94, Lew94]. Algorithm [CI94b, CI95b, DI89, EBSS^{+}11, IK05, CI95a, IK06]. Algorithms [DI90, GD88, Ips87, Ips88, Ips90, Sch88, Ips91]. Analysis [CI95a, CI95b, IK05, IZ20, IK06, IW06, Ips09e]. Angeles [Rod89]. Angle [IM95]. Application [EBSS^{+}11, HI82, HIS18]. Applications [xJ94]. Applied [Lew94, FKL^{+}06]. approach [CI22]. Approximation [DIKMI18, HI15]. Approximations [DI19]. Architectures [DI87a, IS85, IS86, Sch88]. arclength [DKIK07]. array [DI86b]. Arrays [BI87, DI87a, Ips84b]. Attention [IS08, IS07]. Austria [CW86]. aware [HI23].

Backward [CI94a]. Bareiss [DI89, Ips90, Ips91]. based [CI22]. BayesCG [RICO23]. Bayesian [COIG19]. behind [IM98]. Belgium [GD88]. Between [IM95]. Block [DIKMI18]. Bound [HIS18]. Bounds [EI94, EI98b, HI15, CI21, EI99, HI23, Ips00b, Ips00a, Ips03, IR08, IN09].

California [Rod89]. Can [BDI24]. Carlo [EBSS^{+}11, HIS23]. Chains [IM94]. Characteristic [RI11, IR08]. Chicago [DMSV90]. China [IMŠ^{+}09b]. Code [GI89a]. Coefficients [IS11]. Coherence [IW14]. Columns [IW14]. Communicating [FKL^{+}06]. Complementary [IM95]. Complex [Ips96a, Ips00c]. Complexity [ISS84, ISS86]. Components [CI95c]. computability [DI86b]. Computation [DI89, DI90, HIW15, HI15, IS08, IS07]. Computations [Ips83]. Computer [Sch88]. Computing [CI94b, CI95b, DI87b, DMSV90, Ips97, RI11, CI95a, Rod89]. Condition [DKIK07]. Conditioning [HIW15]. cones [DI86b]. Conference [DMSV90, IMŠ^{+}09a, xJ94, Lew94, Rod89, IMŠ^{+}09b]. Conjugate [COIG19]. Conquer [CI94b]. Construction [DI87a]. continuation [DKIK07]. Contributions [Ips96a]. Convergence [CIK^{+}95, CIK^{+}96, DIKMI18, IK05, IK06]. Correlations [DI87b, DI89].

Dangling [IS08, IS07]. Data [DI87b]. December [DMSV90, Rod89]. Decomposition [CI94c, HIW15, Ips84b]. Decompositions [CI94a, HI83]. Deficient [IKP11]. Defined [IS11]. Definite [DI86a, Ips87, Ips88]. Dense [ISS84, ISS86, Ips87, Ips88]. Dense-Linear-System [ISS86]. Dependences [DI90]. determinant [SAI17]. determinants [IR08]. deterministic [HI23]. devoted [IMŠ^{+}09b]. Diagonal [HIS23]. diagonalisable [EI98a]. Digital [GD88]. Discussion [COIG19]. Divide [CI94b]. Do [DI19]. Dominant [DIKMI18]. edition [GIO^{+}19]. Editorial [GIO^{+}19]. Effect [IW14]. Efficient [DI87a]. eigenproblem [Ips10e]. Eigenspaces [EI94]. Eigenvalue [CI94a, IS85, IS86, IJ87a, IJ90, Ips96a, CW86, IJ87b, Ips06e, Ips06f, Ips09f]. Eigenvalues [EI98b, RI11, BI03, EI98a, EI99, Ips98, IN09]. Eigenvector [Ips97]. eigenvectors [EI98a]. Embed [BI85]. Equations [CIK^{+}95, CI95c, CIK^{+}96]. Equivalence [HI82]. Ergodicity [IS11]. Error [IZ20, HI23]. Errors [CI94a]. Estimates [CIK^{+}95, CIK^{+}96, DKIK07]. Estimating [HIS23]. estimators [SAI17]. Europe [CW86]. Exact [HI15]. examples [FKL^{+}06]. excess [CI22]. Expository [Ips10a, Ips10b, Ips10c, Ips10d, Ips11a, Ips11b, Ips11c, Ips11d, Ips12a]. Expressions [Ips00a].

Factorisations [CI91, CI94d]. Fast [Ips84a]. Fifth [Lew94]. floating [HI23]. four [FKL^{+}06]. Fourth [DMSV90]. free [SAI17]. Functions [BKS18].

Gap [DI19]. Generalised [Ips90, Ips91]. Germany [Ips06f]. Givens’ [Ips84a, BI87]. GMRES [CIKM94, CIK^{+}95, CIK^{+}96, CIKM96, Ips00a]. Google [IW06, WI09, WI09]. Gradient [COIG19]. Gram [HI15]. Gray [GI89a].

Hagen [Ips06f]. Hand [BCCI98]. held [CW86, Rod89]. Helmut [Ips96a]. Hermitian [IN09]. History [Ips94, Ips96b]. Hyperbolic [DI86a]. Hypercube [IJ87a, IJ90, IJ87b]. Hypercubes [BI85, GI88, GI89a, GI89b].

IBM [CW86]. idea [IM98]. ILAS

[IMŠ^{+}09a, IMŠ^{+}09b]. Illinois [DMSV90]. Illustration [DI87a]. Impact [IS85, IS86]. Imply [EI98b, EI99]. Importance [EBSS^{+}11]. Improving [JI92]. Inclusion [BI03]. Increase [BDI24]. Influence [BCCI98]. Information [EBSS^{+}11]. Inner [IZ20]. Institute [CW86, GD88]. Integral [CIK^{+}95, CIK^{+}96]. International [IMŠ^{+}09b, xJ94, Ips06f]. Introduction [Ips07a, Ips08a, Ips08d, Ips08e, Ips09d]. invariant [Ips00b, Ips00c, Ips10e]. Inverse [Ips94, Ips96b, Ips97, JI92]. issue [Ips06e, Ips06f, Ips09f, IMŠ^{+}09b]. Iteration [Ips94, Ips96b, Ips97, JI92]. Iterative [CIOR21].

July [CW86, Ips06f, IMŠ^{+}09b]. June [Ips06f, Lew94].

Krylov [DIKMI18, IM98, RICO23].

Lack [BCCI98]. Langville [IK05, IK06]. Large [CW86]. Least [BI87, IKP11, IW14]. Leuven [GD88]. Leverage [HIW15]. Levinson [Ips90, Ips91]. Linear [BCCI98, BI87, CI95c, CIOR21, GD88, ISS84, ISS86, IMŠ^{+}09b, Lew94, BCIH19, CI21, CI22]. Local [GI89a]. log [SAI17]. log-determinant [SAI17]. Low [BKS18, DI19]. Low-Rank [BKS18, DI19]. Lower [BDI24].

Markov [IM94]. Mathematical [IW06]. mathematics [FKL^{+}06]. Matrices [CI94c, IW14, EI98a, Ips00b, Ips00c, Ips01, IN09]. Matrix [BKS18, DI87b, DI19, EI98b, EBSS^{+}11, HIS23, HI15, Ips83, BI03, EI99, Ips98, Ips09e, SAI17]. matrix-free [SAI17]. Mesh [GI88, GI89a, GI89b]. Method [COIG19, Ips84a]. Methodology [DI87a]. Methods [CIOR21, HIS23, IJ87b, IM98]. Meyer [IK05, IK06]. Minimal [CIKM94, CIKM96]. Modern [Sch88]. Monte [EBSS^{+}11, HIS23]. multiple [CI21]. Multiplication [EBSS^{+}11]. Multiplicative [CI21]. Multiprocessor [ISS84, ISS86]. multivariate [CI21].

NATO [GD88]. Need [DI19]. Network [HI82]. Networks [HI83]. Nodes [IS08, IS07]. non [IN09]. non-Hermitian [IN09]. Nonlinear [IKP11]. nonsymmetric [Ips01]. Norms [IS11, CI21]. note [Ips01, Ips03]. Numerical [Ips96a, Ips09e, xJ94, Sch88]. numerics [GIO^{+}19].

Oberlech [CW86]. Ordinal [WI09, WI09]. Orthogonal [HI82, HI83]. Orthonormal [IW14]. overview [Ips00c].

PageRank [IK05, IK06, IW06, IS07, IS08, WI09, WI09]. Papers [Ips10a, Ips10b, Ips10c, Ips10d, Ips11a, Ips11b, Ips11c, Ips11d, Ips12a]. Parallel [DI86a, DI90, DMSV90, GD88, Ips84a, IS85, IS86, Ips87, Ips88, Rod89, Sch88]. Partial [DI87b, DI89]. Perturbation [EI93, EI94, EI95, EI98b, IR08, CI21, EI98a, EI99, Ips98, Ips00b, Ips03, IN09, Ips10e]. point [HI23]. Polynomial [CIKM94, CIKM96]. Polynomials [RI11, IR08]. Positive [DI86a, Ips87, Ips88]. Positive-Definite [Ips87, Ips88]. Precision [BDI24, HI23]. Precision-aware [HI23]. Preconditioned [IW14]. preconditioning [Ips01]. Preface [BDIM15, IMŠ^{+}09a]. prior [RICO23]. Probabilistic [BCIH19, CIOR21, HI15, HIS18, IZ20, GIO^{+}19, HI23]. Problem [IJ87a, IJ90, IJ87b]. Problems [BI87, EI93, EI95, IS85, IS86, Ips96a, Ips05, Ips06a, Ips06b, Ips06c, Ips06d, Ips07b, Ips07c, Ips07d, Ips07e, Ips07f, Ips08d, Ips08e, Ips08b, Ips08c, Ips09d, Ips09a, Ips09b, Ips09c, IKP11, IW14, CW86, Ips06e, Ips06f, Ips09f]. Proceedings [DMSV90, Ips06f, IMŠ^{+}09a, xJ94, Lew94, Rod89, CW86]. Processing [DMSV90, GD88, Rod89]. Products [IZ20]. projector [CI22]. projector-based [CI22]. properties [IW06, RICO23]. pseudo [DKIK07]. pseudo-arclength [DKIK07].

QR [CI95b]. quantifying [CI22].

Randomized [HI15, SAI17]. Rank [BKS18, CI91, CI94d, DI19, IKP11]. Rank-Deficient [IKP11]. Rank-Revealing [CI91, CI94d]. Ranking [WI09, WI09]. Real [HIS23]. Recursive [GI88, GI89b]. Refined [IN09]. Refinement [GI88, GI89a, GI89b]. regions [BI03]. regression [CI21, CI22]. Relative [EI93, EI94, EI95, EI98a, EI98b, Ips98, EI99, Ips00b, Ips00c, Ips03]. Remarks [Ips90, Ips91]. Research [Ips10a, Ips10b, Ips10c, Ips10d, Ips11a, Ips11b, Ips11c, Ips11d, Ips12a, Ips12b]. residual [Ips00a]. Results [DIKMI18, EI98a, Ips98]. Retrieval [EBSS^{+}11]. Revealing [CI91, CI94d]. Right [BCCI98]. Right-Hand [BCCI98]. Ring [ISS84, ISS86]. Rotations [BI87, DI86a, Ips84a].

Sampling [EBSS^{+}11, IW14]. scale [CW86]. Scaled [BI87]. Schatten [CI21]. Scheme [GI89a]. Scientific [DMSV90, Rod89]. Scores [HIW15]. Selection [IKP11]. Sensitivity [CI95c]. Shanghai [IMŠ^{+}09a, IMŠ^{+}09b, xJ94]. SIAM [DMSV90, Lew94, Rod89]. Side [BCCI98]. Signal [GD88]. Singular [BDI24, CI94a, CI94b, CI94c, CI95b, DI19, EI93, EI94, EI95, Ips84b, CI95a, Ips98]. sketched [CI22]. Small [BDI24]. Snowbird [Lew94]. Society [IMŠ^{+}09b]. Solution [BCCI98, BI87, CIK^{+}95, CI95c, DI86a, DI87a, ISS84, IS85, ISS86, IS86, Ips87, Ips88, Ips96a, CIK^{+}96, Ips06e, Ips06f, Ips09f]. solvers [BCIH19]. Solving [IJ87a, IJ90, IJ87b]. Some [Ips90, Ips91]. Spaces [DIKMI18]. Special [Ips06e, Ips06f, IS08, Ips09f, IMŠ^{+}09b, GIO^{+}19, IS07]. Spotlights [Ips12b]. Squares [BI87, IKP11, IW14]. Stability [IM94]. Stable [DI89, Ips83]. Statistical [RICO23]. Structural [DIKMI18]. Study [GD88].

Subset [IKP11]. Subspace [HIS18].

Subspaces [DIKMI18, EI94, HIS18, IM95, Ips00b, Ips00c, Ips10e]. summation [HI23]. Symmetric [DI86a, HIS23, Ips87, IJ87a, Ips88, IJ90, IJ87b]. synthesis [DI86b]. System [BCCI98, ISS84, ISS86]. Systems [CI95c, CIOR21, DI86a, DI87a, Ips87, Ips88]. Systolic [BI87, DI86b, DI87a, HI82, HI83, Ips84b, Ips87, Ips88].

Techniques [EI93, EI95, Ips05, Ips06a, Ips06b, Ips06c, Ips06d, Ips07c, Ips07d, Ips07e, Ips07f, Ips08b, Ips08c, Ips09d, Ips09a, Ips09b, Ips09c, Ips07b, Ips08d, Ips08e]. Their [HI82]. theorems [Ips00c]. theory [Ips10e]. Third [Rod89]. Three [EI98b, EI99]. time [DI86b]. Toeplitz [DI87a, Ips87, Ips88]. total [CI22]. trace [SAI17]. Transformations [HI82]. Trees [BI85]. Triangular [CI94c]. Tridiagonal [IJ87a, IJ90, IJ87b]. Two [IJ87b].

uncertainties [CI22]. Uniform [DI90, GI89a, IM94]. unifying [BCIH19, Ips03]. Untitled [Ips07g, Ips12c]. Updates [BKS18]. Updating [IK05, IK06]. Using [Ips84a]. Utah [Lew94].

Value

[CI94a, CI94c, DI19, EI93, EI95, Ips84b]. Values [BDI24, CI94b, CI95b, CI95a, Ips98]. Vector [EI94, IS11]. view [BCIH19]. VLSI [DI87a, Ips83].

Wielandt [Ips96a]. Workshop [CW86, Ips06f].

References

Banoczi:1998:LIR

[BCCI98] J. M. Banoczi, Nan-Chieh Chiu, Grace E. Cho, and Ilse C. F. Ipsen. The lack of influence of the righthand side on the accuracy of linear system solution. SIAM Journal on Scientific Computing, 20 (1):203–227, January 1998. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30526.

Bartels:2019:PLS

[BCIH19] Simon Bartels, Jon Cockayne, Ilse C. F. Ipsen, and Philipp Hennig. Probabilistic linear solvers: a unifying view. Statistics and Computing, 29(6):1249– 1263, November 2019. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL http://link.springer.
com/article/10.1007/s11222-
019-09897-7;   http://link.
springer.com/content/pdf/10.
1007/s11222-019-09897-7.pdf.

Boutsikas:2024:SSV

[BDI24] Christos Boutsikas, Petros Drineas, and Ilse C. F. Ipsen. Small singular values can increase in lower precision. SIAM Journal on Matrix Analysis and Applications, 45(3):??, ???? 2024. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL https://epubs.siam.org/
doi/10.1137/23M1557209.

Barlow:2015:P

[BDIM15] Jesse L. Barlow, Zlatko Drmac, Ilse Ipsen, and Julio Moro. Preface. Linear Algebra and its Applications, 464(??):1–2, January 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0024379514006533.

Bhatt:1985:HET

[BI85] S. N. Bhatt and I. C. F. Ipsen. How to embed trees in hypercubes. Research Report YALEU/DCS/RR-443, Department of Computer Science, Yale University, 1985.

Barlow:1987:SGR

[BI87] Jesse L. Barlow and Ilse C. F. Ipsen. Scaled Givens rotations for the solution of linear least squares problems on systolic arrays. SIAM Journal on Scientific and Statistical Computing, 8(5): 716–733, 1987. CODEN SIJCD4. ISSN 0196-5204.

Beattie:2003:IRM

[BI03] Christopher Beattie and Ilse C. F. Ipsen. Inclusion regions for matrix eigenvalues. Linear Algebra and its Applications, 358:281– 291, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue on accurate solution of eigenvalue problems (Hagen, 2000).

Beckermann:2018:LRU

[BKS18] Bernhard Beckermann, Daniel Kressner, and Marcel Schweitzer. Low-rank updates of matrix functions. SIAM Journal on Matrix Analysis and Applications, 39(1): 539–565, ???? 2018. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chandrasekaran:1991:RRF

[CI91] S. Chandrasekaran and I. Ipsen.

On rank-revealing QR factorisations. Research Report YALEU/DCS/RR-880, Department of Computer Science, Yale University, New Haven, CT, USA, December 1991. ???? pp.

Chandrasekaran:1994:BEE

[CI94a] S. Chandrasekaran and I. C. F.

Ipsen. Backward errors for eigenvalue and singular value decompositions. Numerische Mathematik, 68(2):215–223, 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandrasekaran:1994:DCA

[CI94b] S. Chandrasekaran and I. C. F.

Ipsen. A divide and conquer algorithm for computing singular values. Z. Angew. Math. Mech., 74 (6):T 532–534, 1994. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Chandrasekaran:1994:SVD

[CI94c] S. Chandrasekaran and I. C. F.

Ipsen. On the singular value decomposition of triangular matrices. In xiong Jiang [xJ94], pages 85–9. ISBN 7-5046-1864-0.

Chandrasekaran:1994:RRF

[CI94d] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. On rank-revealing QR factorisations. SIAM Journal on Matrix Analysis and Applications, 15(2):592–622, 1994. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chandrasekaran:1995:AAC

[CI95a] S. Chandrasekaran and I. C. F. Ipsen. Analysis of a QR algorithm for computing singular values. SIAM Journal on Matrix Analysis and Applications, 16 (2):520–535, 1995. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chandrasekaran:1995:AQA

[CI95b] S. Chandrasekaran and I. C. F. Ipsen. Analysis of a QR algorithm for computing singular values. SIAM Journal on Matrix Analysis and Applications, 16(2): 520–35, 1995. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chandrasekaran:1995:SSC

[CI95c] S. Chandrasekaran and I. C. F. Ipsen. On the sensitivity of solution components in linear systems of equations. SIAM Journal on Matrix Analysis and Applications, 16(1):93–112, 1995. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chi:2021:MPB

[CI21] Jocelyn T. Chi and Ilse C. F. Ipsen. Multiplicative perturbation bounds for multivariate multiple linear regression in Schatten p-norms. Linear Algebra and its Applications, 624(??):87– 102, September 1, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0024379521001518.

Chi:2022:PBA

[CI22] Jocelyn T. Chi and Ilse C. F.

Ipsen. A projector-based approach to quantifying total and excess uncertainties for sketched linear regression. Information and Inference, 11(3):1055–1077, 2022. ISSN 2049-8764.

Campbell:1995:CES

[CIK^{+}95] S. L. Campbell, I. C. F. Ipsen, C. T. Kelley, C. D. Meyer, and Z. Q. Xue. Convergence estimates for solution of integral equations with GMRES. Technical Report CRSC-TR95-12, Center for Research in Scientific Computation, Department of Mathematics, North Carolina State University, 1995.

Campbell:1996:CES

[CIK^{+}96] S. L. Campbell, I. C. F. Ipsen, C. T. Kelley, C. D. Meyer, and Z. Q. Xue. Convergence estimates for solution of integral equations with GMRES. Journal of Integral Equations and Applications, 8 (1):19–34, 1996. ISSN 0897-3962 (print), 1938-2626 (electronic).

Campbell:1994:GMP

[CIKM94] S. L. Campbell, I. C. F. Ipsen, C. T. Kelley, and C. D. Meyer. GMRES and the minimal polynomial. Technical Report CRSC-TR94-10, Center for Research in Scientific Computation, Department of Mathematics, North Carolina State University, 1994.

Campbell:1996:GMP

[CIKM96] S. L. Campbell, I. C. F. Ipsen, C. T. Kelley, and C. D. Meyer. GMRES and the minimal polynomial. BIT Numerical Mathematics, 36(4):664–675, December 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/
BIT/contents/bit36.html.

Cockayne:2021:PIM

[CIOR21] Jon Cockayne, Ilse C. F. Ipsen, Chris J. Oates, and Tim W. Reid. Probabilistic iterative methods for linear systems. Journal of Machine Learning Research (JMLR), 22:Paper No. 232, 34, 2021. CO-DEN JMLRAJ. ISSN 1532-4435 (print), 1533-7928 (electronic).

Cockayne:2019:BCG

[COIG19] Jon Cockayne, Chris J. Oates, Ilse C. F. Ipsen, and Mark Girolami. A Bayesian conjugate gradient method (with discussion). Bayesian Analysis, 14(3):937–1012, September 2019. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL https://

projecteuclid.org/journals/
bayesian-analysis/volume-14/
issue-3/A-Bayesian-Conjugate-
Gradient-Method-with-Discussion/
10.1214/19-BA1145.full. In-

cludes 6 discussions and a rejoinder from the authors.

Cullum:1986:LSE

[CW86] Jane Cullum and Ralph A. Willoughby, editors. Large scale eigenvalue problems: proceedings of the IBM Europe Institute Workshop on Large Scale Eigenvalue Problems held in Oberlech, Austria, July 8–12, 1985. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-70074-9. LCCN QA193 .I261 1985.

Delosme:1986:PSS

[DI86a] Jean-Marc Delosme and Ilse C. F.

Ipsen. Parallel solution of symmetric positive definite systems with hyperbolic rotations. Linear Algebra and its Applications, 77: 75–111, 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special volume on parallel computing.

Delosme:1986:SAS

[DI86b] Jean-Marc Delosme and Ilse C. F.

Ipsen. Systolic array synthesis: computability and time cones. In Parallel algorithms & architectures (Luminy, 1986), pages 295– 312. North-Holland, Amsterdam; New York, 1986.

Delosme:1987:ESA

[DI87a] J.-M. Delosme and I. Ipsen. Efficient systolic arrays for the solution of Toeplitz systems: An illustration of a methodology for the construction of systolic architectures in VLSI. In Systolic Arrays, pages 37–46. Adam Hilger Ltd., Bristol, UK, 1987.

Delosme:1987:CPC

[DI87b] J.-M. Delosme and I. C. F.

Ipsen. Computing partial correlations from the data matrix. Research Report 541, Department of Computer Science, Yale University, 1987.

Delosme:1989:BAS

[DI89] Jean-Marc Delosme and Ilse C. F. Ipsen. From Bareiss’ algorithm to the stable computation of partial correlations. Journal of Computational and Applied Mathematics, 27(1–2):53–91, 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Reprinted in ıt Parallel algorithms for numerical linear algebra, 53–91, North-Holland, Amsterdam, 1990.

Delosme:1990:PCA

[DI90] J.-M. Delosme and I. C. F. Ipsen. Parallel computation of algorithms with uniform dependences. In Dongarra et al. [DMSV90], pages 319–326. ISBN 0-89871-262-9. LCCN QA76.58.S55 1989.

Drineas:2019:LRM

[DI19] Petros Drineas and Ilse C. F. Ipsen. Low-rank matrix approximations do not need a singular value gap. SIAM Journal on Matrix Analysis and Applications, 40 (1):299–319, ???? 2019. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Drineas:2018:SCR

[DIKMI18] Petros Drineas, Ilse C. F. Ipsen, Eugenia-Maria Kontopoulou, and Malik Magdon-Ismail. Structural convergence results for approximation of dominant subspaces from block Krylov spaces.

SIAM Journal on Matrix Analysis and Applications, 39(2):567–586, ???? 2018. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Dickson:2007:CEP

[DKIK07] K. I. Dickson, C. T. Kelley, I. C. F.

Ipsen, and I. G. Kevrekidis. Condition estimates for pseudoarclength continuation. SIAM Journal on Numerical Analysis, 45(1):263–276, 2007. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dongarra:1990:PPS

[DMSV90] Jack Dongarra, Paul Messina, Danny C. Sorensen, and Robert G. Voigt, editors. Parallel Processing for Scientific Computing: Proceedings of the Fourth SIAM Conference on Parallel Processing for Scientific Computing, Chicago, Illinois, December 11–13, 1989. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-262-9. LCCN QA76.58.S55 1989.

Eriksson-Bique:2011:ISM

[EBSS^{+}11] Sylvester Eriksson-Bique, Mary Solbrig, Michael Stefanelli, Sarah Warkentin, Ralph Abbey, and Ilse C. F. Ipsen. Importance sampling for a Monte Carlo matrix multiplication algorithm, with application to information retrieval. SIAM Journal on Scientific Computing, 33(4):1689–1706, ???? 2011. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/

                                        sisc/resource/1/sjoce3/v33/
                                        i4/p1689_s1.

Eisenstat:1993:RPT

[EI93] Stanley C. Eisenstat and Ilse C. F. Ipsen. Relative perturbation techniques for singular value problems. Research Report YALEU/DCS/RR-942, Department of Computer Science, Yale University, New Haven, CT, USA, July 1993. ???? pp.

Eisenstat:1994:RPB

[EI94] S. C. Eisenstat and I. C. F. Ipsen. Relative perturbation bounds for eigenspaces and singular vector subspaces. In Lewis [Lew94], pages 62–65. ISBN 0-89871-336-6. LCCN QA184.S58 1994.

Eisenstat:1995:RPT

[EI95] Stanley C. Eisenstat and Ilse C. F. Ipsen. Relative perturbation techniques for singular value problems. SIAM Journal on Numerical Analysis, 32(6):1972–1988, December 1995. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eisenstat:1998:RPR

[EI98a] S. C. Eisenstat and I. C. F. Ipsen. Relative perturbation results for eigenvalues and eigenvectors of diagonalisable matrices. BIT, 38(3): 502–509, 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Eisenstat:1998:TAP

[EI98b] Stanley C. Eisenstat and Ilse C. F. Ipsen. Three absolute perturbation bounds for matrix eigenvalues imply relative bounds. SIAM Journal on Matrix Analysis and Applications, 20(1):149– 158, 1998. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/
article/32328.

Eisenstat:1999:TAP

[EI99] Stanley C. Eisenstat and Ilse C. F. Ipsen. Three absolute perturbation bounds for matrix eigenvalues imply relative bounds. SIAM Journal on Matrix Analysis and Applications, 20(1):149– 158, 1999. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Finkel:2006:CAM

[FKL^{+}06] Daniel E. Finkel, Christopher Kuster, Matthew Lasater, Rachel Levy, Jill P. Reese, and Ilse C. F. Ipsen. Communicating applied mathematics: four examples. SIAM Review, 48(2):359– 389, 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golub:1988:NAS

[GD88] Gene H. Golub and Paul Van Dooren, editors. NATO Advanced Study Institute on Linear Algebra, Digital Signal Processing, and Parallel Algorithms (1988: Leuven, Belgium), volume 70 of NATO ASI series. Series F, Computer and systems sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-52300-6 (New York), 3-540-52300-6 (Berlin). LCCN QA184 .N39 1988.

Gropp:1988:RMR

[GI88] W. Gropp and I. Ipsen. Recursive mesh refinement on hypercubes. Technical Report RR-616, Department of Computer Science, Yale University, New Haven, CT, USA, 1988.

Gropp:1989:GCS

[GI89a] W. D. Gropp and I. C. F. Ipsen. A Gray code scheme for local uniform mesh refinement on hypercubes. In Rodrigue [Rod89], pages 202–6. ISBN 0-89871-228-9. LCCN QA76.5 .C61921 1987.

Gropp:1989:RMR

[GI89b] W. D. Gropp and I. C. F. Ipsen. Recursive mesh refinement on hypercubes. BIT, 29(2):186– 211, 1989. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Girolami:2019:ESE

[GIO^{+}19] M. Girolami, I. C. F. Ipsen, C. J. Oates, A. B. Owen, and T. J. Sullivan. Editorial: special edition on probabilistic numerics. Statistics and Computing, 29(6): 1181–1183, November 2019. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.
                                        com/article/10.1007/s11222-
                                        019-09892-y;   http://link.
                                        springer.com/content/pdf/10.
                                        1007/s11222-019-09892-y.pdf.

Heller:1982:SNO

[HI82] D. Heller and I. Ipsen. Systolic network for orthogonal equivalence transformations and their application. In Proceedings of the Conference on Advanced Research in VLSI, pages 113–122. MIT Press, Cambridge, MA, USA, 1982.

Heller:1983:SNO

[HI83] Don E. Heller and Ilse C. F. Ipsen.

Systolic networks for orthogonal decompositions. SIAM Journal on Scientific and Statistical Computing, 4(2):261–269, 1983. CODEN SIJCD4. ISSN 0196-5204.

Holodnak:2015:RAG

[HI15] John T. Holodnak and Ilse C. F.

Ipsen. Randomized approximation of the Gram matrix: Exact computation and probabilistic bounds. SIAM Journal on Matrix Analysis and Applications, 36(1): 110–137, ???? 2015. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Hallman:2023:PAD

[HI23] Eric Hallman and Ilse C. F.

Ipsen. Precision-aware deterministic and probabilistic error bounds for floating point summation. Numerische Mathematik, 155(1–2): 83–119, October 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.
com/article/10.1007/s00211-
023-01370-y.

Holodnak:2018:PSB

[HIS18] John T. Holodnak, Ilse C. F. Ipsen, and Ralph C. Smith. A probabilistic subspace bound with application to active subspaces. SIAM Journal on Matrix Analysis and Applications, 39(3):1208– 1220, ???? 2018. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Hallman:2023:MCM

[HIS23] Eric Hallman, Ilse C. F. Ipsen, and Arvind K. Saibaba. Monte Carlo methods for estimating the diagonal of a real symmetric matrix. SIAM Journal on Matrix Analysis and Applications, 44(1):??, ???? 2023. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL https://epubs.siam.org/
doi//doi/10.1137/22M1476277.

Holodnak:2015:CLS

[HIW15] John T. Holodnak, Ilse C. F. Ipsen, and Thomas Wentworth. Conditioning of leverage scores and computation by QR decomposition. SIAM Journal on Matrix Analysis and Applications, 36 (3):1143–1163, ???? 2015. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:1987:SST

[IJ87a] I. Ipsen and E. Jessup. Solving the symmetric tridiagonal eigenvalue problem on the hypercube. Technical Report YALEU/DCS/RR-548, Department of Computer Science, Yale University, New Haven, CT, USA, 1987.

Ipsen:1987:TMS

[IJ87b] Ilse C. F. Ipsen and Elizabeth R.

Jessup. Two methods for solving the symmetric tridiagonal eigenvalue problem on the hypercube. In Hypercube multiprocessors 1987 (Knoxville, TN, 1986), pages 627– 638. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987.

Ipsen:1990:SST

[IJ90] Ilse C. F. Ipsen and Elizabeth R.

Jessup. Solving the symmetric tridiagonal eigenvalue problem on the hypercube. SIAM Journal on Scientific and Statistical Computing, 11(2):203–229, 1990. CODEN SIJCD4. ISSN 0196-5204.

Ipsen:2005:CAP

[IK05] Ilse C. F. Ipsen and Steve Kirkland. Convergence analysis of a PageRank updating algorithm by Langville and Meyer. SIAM Journal on Matrix Analysis and Applications, 27(4):952–967, 2005. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2006:CAP

[IK06] Ilse C. F. Ipsen and Steve Kirkland. Convergence analysis of a PageRank updating algorithm by Langville and Meyer. SIAM Journal on Matrix Analysis and Applications, 27(4):952–967, 2006. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2011:RDN

[IKP11] I. C. F. Ipsen, C. T. Kelley, and S. R. Pope. Rank-deficient nonlinear least squares problems and subset selection. SIAM Journal on Numerical Analysis, 49 (3):1244–1266, ???? 2011. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

URL http://epubs.siam.org/
sinum/resource/1/sjnaam/v49/
i3/p1244_s1.

Ipsen:1994:USM

[IM94] Ilse C. F. Ipsen and Carl D. Meyer. Uniform stability of Markov chains. SIAM Journal on Matrix Analysis and Applications, 15(4):1061–1074, 1994. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:1995:ABC

[IM95] Ilse C. F. Ipsen and Carl D. Meyer. The angle between complementary subspaces. American Mathematical Monthly, 102(10): 904–911, ???? 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ipsen:1998:IBK

[IM98] Ilse C. F. Ipsen and Carl D. Meyer. The idea behind Krylov methods. American Mathematical Monthly, 105(10):889–899, 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ipsen:2009:PIC

[IMŠ^{+}09a] Ilse Ipsen, Julio Moro, Peter Šemrl, Jia-Yu Shao, and Pei Yuan Wu. Preface to the 14th ILAS Conference Proceedings, Shanghai 2007. Linear Algebra and its Applications, 430(5–6):1441, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Held at Shanghai University, Shanghai, July 16–20, 2007.

Ipsen:2009:SID

[IMŠ^{+}09b] Ilse Ipsen, Julio Moro, Peter Šemrl, Jia-Yu Shao, and Pei Yuan Wu. Special issue devoted to the 14th ILAS (International Linear Algebra Society) conference, Shanghai, China, July 16– 20, 2007. Linear Algebra and its Applications, 430(5–6):1441– 1724, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ipsen:2009:RPB

[IN09] I. C. F. Ipsen and B. Nadler.

Refined perturbation bounds for eigenvalues of Hermitian and non-Hermitian matrices. SIAM Journal on Matrix Analysis and Applications, 31(1):40–53, 2009. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:1983:SMC

[Ips83] Ilse C. F. Ipsen. Stable Matrix Computations in VLSI. Ph.D. thesis, Department of Computer Science, The Pennsylvania State University, University Park, PA, USA, 1983. 198 pp.

Ipsen:1984:PMU

[Ips84a] I. Ipsen. A parallel QR method using fast Givens’ rotations. Technical Report YALEU/DCS/RR-299, Department of Computer Science, Yale University, New Haven, CT, USA, 1984.

Ipsen:1984:SVD

[Ips84b] I. Ipsen. Singular value decomposition with systolic arrays. In Proceedings of the Society of Photo-Optical Engineers. Society of Photo-Optical Engineers, Bellingham, WA, 1984.

Ipsen:1987:SAP

[Ips87] I. Ipsen. Systolic algorithms for the parallel solution of dense symmetric positive-definite Toeplitz systems. Technical Report YALEU/DCS/RR-539, Department of Computer Science, Yale University, New Haven, CT, USA, May 1987.

Ipsen:1988:SAP

[Ips88] Ilse C. F. Ipsen. Systolic algorithms for the parallel solution of dense symmetric positive-definite Toeplitz systems. In Schultz [Sch88], pages 85–108. ISBN 0-387-96733-8. LCCN QA76.5.N79 1988.

Ipsen:1990:SRG

[Ips90] I. C. F. Ipsen. Some remarks on the generalised Bareiss and Levinson algorithms. In Golub and Dooren [GD88], pages 189–214. ISBN 0-387-52300-6 (New York), 3-540-52300-6 (Berlin). LCCN QA184 .N39 1988.

Ipsen:1991:SRG

[Ips91] Ilse Ipsen. Some remarks on the generalised Bareiss and Levinson algorithms. In Numerical linear algebra, digital signal processing and parallel algorithms (Leuven, 1988), volume 70 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 189–214. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Ipsen:1994:HII

[Ips94] Ilse C. F. Ipsen. A history of inverse iteration. In B. Huppert and H. Schneider, editors, Helmut Wielandt, Mathematische Werke, Mathematical Works, Volume II: Matrix Theory and Analysis, page ?? Walter de Gruyter, Berlin, Germany, 1994. To appear.

Ipsen:1996:HWC

[Ips96a] I. C. F. Ipsen. Helmut Wielandt’s contributions to the numerical solution of complex eigenvalue problems. In Bertram Huppert and Hans Schneider, editors, Helmut Wielandt, Mathematische Werke, Mathematical Works, volume II: Matrix Theory and Analysis, pages 453–463. Walter de Gruyter, Berlin, Germany, 1996. ISBN 3-11-012452-1. LCCN QA3.W52 1993.

Ipsen:1996:HII

[Ips96b] I. C. F. Ipsen. A history of inverse iteration. In Bertram Huppert and Hans Schneider, editors, Helmut Wielandt, Mathematische Werke, Mathematical Works, volume II: Matrix Theory and Analysis, pages 464–472. Walter de Gruyter, Berlin, Germany, 1996. ISBN 3-11-012452-1. LCCN QA3.W52 1993.

Ipsen:1997:CEI

[Ips97] Ilse C. F. Ipsen. Computing an eigenvector with inverse iteration. SIAM Review, 39(2):254–291, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/dbq/
                                        article/30077.

Ipsen:1998:RPR

[Ips98] Ilse C. F. Ipsen. Relative perturbation results for matrix eigenvalues and singular values. Acta Numerica, 7:151–201, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Ipsen:2000:EBG

[Ips00a] I. C. F. Ipsen. Expressions and bounds for the GMRES residual. BIT Numerical Mathematics, 40(3):524–535, September 2000. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Ipsen:2000:ARP

[Ips00b] Ilse C. F. Ipsen. Absolute and relative perturbation bounds for invariant subspaces of matrices. Linear Algebra and its Applications, 309(1–3):45– 56, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

                                        URL http://www.elsevier.nl/
                                        gej-ng/10/30/19/126/25/28/
                                        abstract.html;   http://www.

elsevier.nl/gej-ng/10/30/19/
126/25/28/article.pdf. Pro-

ceedings of the International Workshop on Accurate Solution of Eigenvalue Problems (University Park, PA, 1998).

Ipsen:2000:ORT

[Ips00c] Ilse C. F. Ipsen. An overview of relative sin Θ theorems for invariant subspaces of complex matrices. Journal of Computational and Applied Mathematics, 123 (1–2):131–153, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis 2000, Vol. III. Linear algebra.

Ipsen:2001:NPN

[Ips01] Ilse C. F. Ipsen. A note on preconditioning nonsymmetric matrices. SIAM Journal on Scientific Computing, 23(3):1050– 1051, 2001. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Ipsen:2003:NUA

[Ips03] Ilse C. F. Ipsen. A note on unifying absolute and relative perturbation bounds. Linear Algebra and its Applications, 358:239– 253, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue on accurate solution of eigenvalue problems (Hagen, 2000).

Ipsen:2005:PT

[Ips05] Ilse Ipsen. Problems and techniques. SIAM Review, 47(4):

707, December 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97156.

Ipsen:2006:PTa

[Ips06a] Ilse Ipsen. Problems and techniques. SIAM Review, 48(1):41– 42, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:PTb

[Ips06b] Ilse Ipsen. Problems and techniques. SIAM Review, 48(2):305, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:PTc

[Ips06c] Ilse Ipsen. Problems and techniques. SIAM Review, 48(3):485, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:PTd

[Ips06d] Ilse Ipsen. Problems and techniques. SIAM Review, 48(4): 679–680, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:SIA

[Ips06e] Ilse C. F. Ipsen. Special issue on accurate solution of eigenvalue problems. SIAM Journal on Matrix Analysis and Applications, 28 (4):ix, 2006. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2006:SIP

[Ips06f] Ilse C. F. Ipsen. Special issue: Proceedings of the international workshop on accurate solution of eigenvalue problems, Hagen, Germany, June 29–July 1, 2004. SIAM Journal on Matrix Analysis and Applications, 28(4): 907–1203, 2006. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2007:I

[Ips07a] Ilse Ipsen. Introduction. SIAM Review, 49(4):593, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PT

[Ips07b] Ilse Ipsen. Problems and techniques. SIAM Review, 49(1):33– 34, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTa

[Ips07c] Ilse Ipsen. Problems and techniques. SIAM Review, 49(1):33– 34, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTb

[Ips07d] Ilse Ipsen. Problems and techniques. SIAM Review, 49(2):209, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTc

[Ips07e] Ilse Ipsen. Problems and techniques. SIAM Review, 49(3): 419–420, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTd

[Ips07f] Ilse Ipsen. Problems and techniques. SIAM Review, 49(4):593, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:U

[Ips07g] Ilse Ipsen. Untitled. SIAM Review, 49(3):419–420, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:I

[Ips08a] Ilse Ipsen. Introduction. SIAM Review, 50(1):35, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTa

[Ips08b] Ilse Ipsen. Problems and techniques. SIAM Review, 50(1):35, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTd

[Ips08c] Ilse Ipsen. Problems and techniques. SIAM Review, 50(4):721, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTIa

[Ips08d] Ilse Ipsen. Problems and techniques — introduction. SIAM Review, 50(3):483, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTIb

[Ips08e] Ilse Ipsen. Problems and techniques — introduction. SIAM Review, 50(2):273, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTa

[Ips09a] Ilse Ipsen. Problems and techniques. SIAM Review, 51(1):127, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTb

[Ips09b] Ilse Ipsen. Problems and techniques. SIAM Review, 51(2):337, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTc

[Ips09c] Ilse Ipsen. Problems and techniques. SIAM Review, 51(3):543, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTI

[Ips09d] Ilse Ipsen. Problems and techniques: Introduction. SIAM Review, 51(4):705, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:NMA

[Ips09e] Ilse C. F. Ipsen. Numerical matrix analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2009. ISBN 0-89871-676-4. xiv + 128 pp. LCCN QA275 .I67 2009. URL http://www.

loc.gov/catdir/enhancements/
fy0912/2009008514-b.htm;
http://www.loc.gov/catdir/
enhancements/fy0912/2009008514-
d.htm;   http://www.loc.gov/
catdir/enhancements/fy0912/
2009008514-t.htm.

Ipsen:2009:SIA

[Ips09f] Ilse C. F. Ipsen. Special issue on accurate solution of eigenvalue problems. SIAM Journal on Matrix Analysis and Applications, 31 (1):vii, 2009. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2010:ERPa

[Ips10a] Ilse Ipsen. Expository research papers. SIAM Review, 52(1):55, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPb

[Ips10b] Ilse Ipsen. Expository research papers. SIAM Review, 52(2):267, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPc

[Ips10c] Ilse Ipsen. Expository research papers. SIAM Review, 52(3):453, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPd

[Ips10d] Ilse Ipsen. Expository research papers. SIAM Review, 52(4):675, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:EIS

[Ips10e] Ilse C. F. Ipsen. The eigenproblem and invariant subspaces: perturbation theory. In G. W. Stewart: selected works with commentaries, Contemp. Mathematicians, pages 71–93. Birkhäuser Verlag, Basel, Switzerland, 2010.

Ipsen:2011:ERPa

[Ips11a] Ilse Ipsen. Expository research papers. SIAM Review, 53(1):69, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sirev/resource/1/siread/v53/
i1/p69_s1.

Ipsen:2011:ERPb

[Ips11b] Ilse Ipsen. Expository research papers. SIAM Review, 53(2):289, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sirev/resource/1/siread/v53/
i2/p289_s1.

Ipsen:2011:ERPc

[Ips11c] Ilse Ipsen. Expository research papers. SIAM Review, 53(3):503, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sirev/resource/1/siread/v53/
i3/p503_s1.

Ipsen:2011:ERPd

[Ips11d] Ilse Ipsen. Expository research papers. SIAM Review, 53(4):721, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sirev/resource/1/siread/v53/
i4/p721_s1.

Ipsen:2012:ERP

[Ips12a] Ilse Ipsen. Expository research papers. SIAM Review, 53(4):721, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sirev/resource/1/siread/v53/
i4/p721_s1.

Ipsen:2012:RS

[Ips12b] Ilse Ipsen. Research spotlights. SIAM Review, 54(1): 119, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://epubs.siam.org/
sirev/resource/1/siread/v54/
i1/p119_s1.

Ipsen:2012:U

[Ips12c] Ilse Ipsen. Untitled. SIAM Review, 54(1):119–120, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PBD

[IR08] Ilse C. F. Ipsen and Rizwana Rehman. Perturbation bounds for determinants and characteristic polynomials. SIAM Journal on Matrix Analysis and Applications, 30(2):762–776, 2008. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:1985:IPA

[IS85] I. Ipsen and Y. Saad. The impact of parallel architectures on the solution of eigenvalue problems. Technical Report YALEU/DCS/RR-444, Department of Computer Science, Yale University, New Haven, CT, USA, December 1985.

Ipsen:1986:IPA

[IS86] I. C. F. Ipsen and Y. Saad. The impact of parallel architectures on the solution of eigenvalue problems. In Cullum and Willoughby [CW86], pages 37–49. ISBN 0-444-70074-9. LCCN QA193 .I261 1985.

Ipsen:2007:PCS

[IS07] Ilse C. F. Ipsen and Teresa M.

Selee. PageRank computation, with special attention to dangling nodes. SIAM Journal on Matrix Analysis and Applications, 29(4): 1281–1296, 2007. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2008:PCS

[IS08] Ilse C. F. Ipsen and Teresa M.

Selee. PageRank computation, with special attention to dangling nodes. SIAM Journal on Matrix Analysis and Applications, 29(4): 1281–1296, ???? 2008. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2011:ECD

[IS11] Ilse C. F. Ipsen and Teresa M. Selee. Ergodicity coefficients defined by vector norms. SIAM Journal on Matrix Analysis and Applications, 32(1):153–200, ???? 2011. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (elec-

tronic). URL http://epubs.
siam.org/simax/resource/1/
sjmael/v32/i1/p153_s1.

Ipsen:1984:CDL

[ISS84] I. Ipsen, Y. Saad, and M. H. Schultz. Complexity of dense linear system solution on a multiprocessor ring. Technical Report 349, Computer Science Department, Yale University, New Haven, CT, USA, 1984.

Ipsen:1986:CDL

[ISS86] Ilse C. F. Ipsen, Youcef Saad, and Martin H. Schultz. Complexity of dense-linear-system solution on a multiprocessor ring. Linear Algebra and its Applications, 77:205– 239, 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special volume on parallel computing.

Ipsen:2006:MPA

[IW06] Ilse C. F. Ipsen and Rebecca S. Wills. Mathematical properties and analysis of Google’s PageRank. Bol. Soc. Esp. Mat. Apl.

Se⃗MA, 34:191–196, 2006. ISSN 1575-9822.

Ipsen:2014:ECS

[IW14] Ilse C. F. Ipsen and Thomas Wentworth. The effect of coherence on sampling from matrices with orthonormal columns, and preconditioned least squares problems. SIAM Journal on Matrix Analysis and Applications, 35(4): 1490–1520, ???? 2014. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Ipsen:2020:PEA

[IZ20] Ilse C. F. Ipsen and Hua Zhou.

Probabilistic error analysis for inner products. SIAM Journal on Matrix Analysis and Applications, 41(4):1726–1741, ???? 2020. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Jessup:1992:IAI

[JI92] Elizabeth R. Jessup and Ilse C. F.

Ipsen. Improving the accuracy of inverse iteration. SIAM Journal on Scientific and Statistical Computing, 13(2):550–572, 1992. CO-DEN SIJCD4. ISSN 0196-5204.

Lewis:1994:ALA

[Lew94] John G. Lewis, editor. Applied Linear Algebra: Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, Snowbird, Utah, June 15–18, 1994. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-336-6. LCCN QA184.S58 1994.

Rehman:2011:CCP

[RI11] Rizwana Rehman and Ilse C. F. Ipsen. Computing characteristic polynomials from eigenvalues. SIAM Journal on Matrix Analysis and Applications, 32(1):90–114, ???? 2011. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/
simax/resource/1/sjmael/v32/
i1/p90_s1.

Reid:2023:SPB

[RICO23] Tim W. Reid, Ilse C. F. Ipsen, Jon Cockayne, and Chris J. Oates. Statistical properties of BayesCG under the Krylov prior. Numerische Mathematik, 155(3-4): 239–288, December 2023. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.
com/article/10.1007/s00211-
023-01375-7.

Rodrigue:1989:PPS

[Rod89] Garry Rodrigue, editor. Parallel processing for scientific computing: Proceedings of the Third SIAM Conference held in Los Angeles, California, December 1–4, 1987. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1989. ISBN 0-89871-228-9. LCCN QA76.5 .C61921 1987.

Saibaba:2017:RMF

[SAI17] Arvind K. Saibaba, Alen Alexanderian, and Ilse C. F. Ipsen. Randomized matrix-free trace and log-determinant estimators. Numerische Mathematik, 137(2): 353–395, October 2017. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.
com/article/10.1007/s00211-
017-0880-z.

Schultz:1988:NAM

[Sch88] Martin Schultz, editor. Numerical Algorithms for Modern Parallel Computer Architectures, volume 13 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96733-8. vi + 232 pp. LCCN QA76.5.N79 1988.

Wills:2008:ORG

[WI09] Rebecca S. Wills and Ilse C. F.

Ipsen. Ordinal ranking for Google’s PageRank. SIAM Journal on Matrix Analysis and Applications, 30(4):1677–1696, 2008/09. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Wills:2009:ORG

[WI09] Rebecca S. Wills and Ilse C. F.

Ipsen. Ordinal ranking for Google’s PageRank. SIAM Journal on Matrix Analysis and Applications, 30(4):1677–1696, ???? 2009. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Jiang:1994:NAP

[xJ94] Er xiong Jiang, editor. Numerical Algebra, Proceedings of ’92 Shanghai International Numerical Algebra and Its Applications Conference. China Science and Technology Press, Shanghai, China, 1994. ISBN 7-5046-1864-0.